OpenStudy (anonymous):

Which equation shows the formula A = pe^(rt) solved for t ? A. t=In(Ap)-In r B. t=ln (A/p)-ln r C. t= ln(A/p)/r D. t= ln(Ap)/r

OpenStudy (tkhunny):

Why don't you start with the original equation and solve it for 't'?

OpenStudy (anonymous):

plzz help me

OpenStudy (whpalmer4):

Some properties of logs that might help: \[\ln e^x = x\]

OpenStudy (whpalmer4):

\[A=pe^{rt}\]Divide both sides by p \[\frac Ap = e^{rt}\]Take natural log of both sides and you get?

OpenStudy (anonymous):

no clue:/

OpenStudy (whpalmer4):

What's the natural log of the left side?

OpenStudy (whpalmer4):

Why did you close the question?

OpenStudy (tkhunny):

Dear @venom936. If you REALLY have no clue, please try to figure out why you are in this class. If you are being introduced to logarithms, you MUST have some background in simple algebraic manipulation. In the future, please DO some work on your own and SHOW us YOUR work. In this way, YOU will learn something.